Find the seventh term of the geometric sequence with these first three terms.
-5/9, -5/3, -5

Answer by edjones(8007)

-5/3 / -5/9
=-5/3 * -9/5 = 3
r=3
.
a[n]=a[1]r^(n-1)
a[7]=-5/9 * 3^6
=-5/9 * 729
= -405
.
Ed